OpenStudy (johnnydicamillo):

Evaluate the integral x cos5x dx

OpenStudy (johnnydicamillo):

okay so I need to do integration by parts

OpenStudy (johnnydicamillo):

u = x and dv = cos5x

OpenStudy (johnnydicamillo):

then du = 1 and v = -\[\frac{ 1 }{ 5 } \sin 5x\]

OpenStudy (johnnydicamillo):

i think

OpenStudy (johnnydicamillo):

wait no negative

OpenStudy (johnnydicamillo):

we use the forumla \[uv - \int\limits_{}^{}v du\]

OpenStudy (johnnydicamillo):

so I think I get \[\frac{ 1 }{ 5 }x \sin(5x) + \frac{ 1 }{ 25 } \cos(5x) + c\] but I am not sure